OpenStudy (anonymous):

The management of a large store wishes to add a fenced in rectangular storage yard of 20000 ft^2 using the building as one side of the yard. Find the minimum amount of fencing that must be used to enclose the remaining 3 sides of the yard.

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay, I am confused on what formula to use

OpenStudy (anonymous):

Called away a short time ago. area = h*w, 20000= h*w perimeter = 2h + w Solve 20000 = h*w for w and plug the result into the RHS of the perimeter equation. Take the derivative of the resulting RHS, set the derivative expression to zero and solve for h.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome and thank you for the medal. If you have access to the Mathematica 8 program, then this problem could have been solved using the Constrained Optimization function, Minimize: \[\text{Minimize}[\{2h+w,w h==20000,w>0,h>0\},\{w,h\}] \]\[\{400,\{w\to 200,h\to 100\}\} \] The 400 is the minimum perimeter value. No knowledge of the Calculus required with this approach.

OpenStudy (anonymous):

Oh, okay, thanks I didn't know about that! Thank you !